# Non-commutativity as a Universal Characterization for Enhanced Quantum Metrology

### v1

@@ Abstract
A central challenge in quantum metrology is to effectively harness quantum resources to surpass classical precision bounds. Although recent studies suggest that the indefinite causal order may enable sensitivities to attain the super-Heisenberg scaling, the physical origins of such enhancements remain elusive. Here, we introduce the nilpotency index $\mathcal{K}$, which quantifies the depth of non-commutativity between operators during the encoding process, can act as a fundamental parameter governing quantum-enhanced sensing. We show that a finite $\mathcal{K}$ yields an enhanced scaling of root-mean-square error as $N^{-(1+\mathcal{K})}$. Meanwhile, the requirement for indefinite causal order arises only when the nested commutators become constant. Remarkably, in the limit $\mathcal{K}\to\infty$, exponential precision scaling $N^{-1}e^{-N}$ is achievable. We propose experimentally feasible protocols implementing these mechanisms, providing a systematic pathway towards practical quantum-enhanced metrology.
@@
Introduction.—Quantum metrology represents one of the most promising near-term applications of quantum technologies [1, 2, 3, 4, 5, 6]. Its fundamental advantage stems from the ability of quantum resources—such as entanglement and coherence—to enhance measurement precision beyond the limits achievable with classical methods [7, 8, 9]. This is exemplified by the reduction in the root-mean-square error (RMSE) for parameter estimation from the standard quantum limit, scaling as $N^{-1/2}$, to the Heisenberg limit, scaling as $N^{-1}$, when utilizing $N$ resources [10, 11]. Such Heisenberg scaling has been demonstrated across diverse experimental platforms, achieved either by preparing $N$ entangled probes measured in parallel [12, 13, 14] or by employing a single probe that evolves through and is measured after $N$ repetitions [Fig. 1(a)] [15, 16, 17].
@@
To further push the sensitivity limit in measurement precision, recent processes involving indefinite causal order (ICO), where the sequence of fundamental encoding operations lacks definite temporal sequences, have been incorporated into quantum metrology [18, 19, 20, 21, 22, 23]. A paradigmatic example is the quantum SWITCH [Fig. 1(b)], in which a control quantum system coherently dictates the order of two operations applied to a target system [24, 25, 26, 27, 28, 29, 30]. Significantly, using state-of-the-art photonic techniques, a precision scaling with super-Heisenberg limit $N^{-2}$ has been demonstrated [31, 32].
@@
However, the fundamental mechanism by which ICO enhances measurement precision remains poorly understood, representing a key foundational question [33, 32, 34, 35, 36, 37, 38]. Understanding the essential resource underpinning this advantage holds significant promise for reducing experimental overhead while advancing the ultimate sensitivity limits. While it is often assumed that the coherence of causal order plays a central role, evidence suggests that simply increasing the number of orders does not necessarily lead to improved precision [39]. In existing protocols performing displacements generated by a given quadrature $\hat{X}$ and the conjugate quadrature $\hat{P}$ [33, 32, 31] or two unitary gates $\hat{U}$ and $\hat{U}^{\dagger}$ for parameter encoding [34], continuing to add similar paths brings little benefit and even makes the experiment harder to implement. Instead, the enhancement appears to stem from the non-commutativity between encoding operations in two paths, highlighting it as a more critical factor than only order multiplicity.
@@
Here, we establish a quantum metrology framework that exploits non-commutative sequences of encoding operations to enhance measurement precision. Central to our approach is the introduction of the non-commutativity depth $\mathcal{K}$, which is defined as the maximal integer yielding a non-zero nested commutator and serves as a fundamental quantity characterizing the non-commutative structure of the encoding operators. We rigorously demonstrate that $\mathcal{K}$ universally governs the scaling behavior of RMSE. For finite $\mathcal{K}$, the RMSE scales as $N^{-(1+\mathcal{K})}$ when the nested commutator remains non-constant, enabling precision enhancement even without requiring the ICO scheme. However, when the $\mathcal{K}$-th order commutator becomes constant, ICO becomes necessary to unlock the optimal scaling. In the asymptotic limit $\mathcal{K}\to\infty$, we show that an exponential RMSE scaling $N^{-1}e^{-N}$ can be achieved, provided the nested commutators exhibit non-zero variance on the relevant initial state. Furthermore, we propose experimentally feasible protocols capable of achieving these enhanced scalings. Our findings offer a systematic approach to designing quantum control strategies for metrological optimization under realistic experimental conditions.
@@
General framework.—Generally, the precision of quantum metrology estimation governed by a unitary encoding $\hat{U}_{\lambda}=\text{exp}(-i\lambda\hat{H}_{\lambda})$ is fundamentally determined by the distinguishability between the neighboring output states $|\Psi_{\lambda}\rangle=\hat{U}_{\lambda}|\Psi\rangle$ and $|\Psi_{\lambda+\delta\lambda}\rangle=\hat{U}_{\lambda+\delta\lambda}|\Psi\rangle$. This distinguishability is quantified by the quantum Fisher information (QFI) $F_{\lambda}=4\text{Var}[\hat{h}_{\lambda}]_{|\Psi\rangle}$, where $\hat{h}_{\lambda}=i\hat{U}_{\lambda}^{\dagger}\partial_{\lambda}\hat{U}_{\lambda}$ denotes the local generator associated with $\lambda$, and the variance $\text{Var}[\cdots]_{|\Psi\rangle}$ is evaluated to the probe state $|\Psi\rangle$. The QFI sets a fundamental lower bound on the RMSE via the quantum Cramér-Rao bound (QCRB) $\Delta\lambda=1/\sqrt{\nu F_{\lambda}}$, where $\nu$ is the number of independent measurements. This bound is saturable in the asymptotic limit $\nu\to\infty$ through optimal choice of measurement and estimator [40, 41, 42, 43].
@@
Figure 1: The schematic illustrates representative quantum estimation schemes: (a) In the sequential protocol, a single probe state sequentially undergoes NN times of the parameter-encoding operator $\hat{H}_{\lambda}$. (b) In the ICO-based scheme, the probe evolves through a non-commutative encoding sequence whose order is coherently controlled by an auxiliary qubit. (c) In the general non-commutative encoding framework, the probe undergoes a sequence of $2N$ non-commutative encoding operations—NN times $\hat{H}_{g}$ and NN times $\hat{H}_{\lambda}$—executed in a definite order.
@@
We begin by introducing the general framework of the non-commutative metrology scheme [Fig. 1(c)]. The encoding unitary $\hat{U}_{\lambda}$ is constructed by sequentially applying two non-commutative operators, $\hat{H}_{\lambda}$ and $\hat{H}_{g}$ (i.e., $\hat{H}_{g},\hat{H}_{\lambda}]\neq 0$), and takes the form
@@
To quantify the depth of non-commutativity between $\hat{H}_{\lambda}$ and $\hat{H}_{g}$, we define the nilpotency index $\mathcal{K}\in\mathbb{N}$ via the adjoint action $\mathrm{ad}_{\hat{A}}(\hat{B})\equiv[\hat{A},\hat{B}$], such that
@@
We then define $\mathcal{K}$ as the maximal integer for which the nested commutator remains non-zero:
@@
That is, the action of $\mathrm{ad}_{\hat{H}_{g}}$ on $\hat{H}_{\lambda}$ terminates at order $\mathcal{K}+1$, $(\mathrm{ad}_{\hat{H}_{g}})^{\mathcal{K}+1}(\hat{H}_{\lambda})=0,$ which implies that $\mathrm{ad}_{\hat{H}_{g}}$ is nilpotent of index $\mathcal{K}+1$ on $\hat{H}_{\lambda}$. We note that the nilpotency index $\mathcal{K}$ is determined by the non-commutative algebraic structure and is independent of how many times the encoding operators are applied sequentially.
@@
Under the conditions of $\overline{g}\neq 0$ and a non-zero variance $\text{Var}[[\hat{H}_{g},\hat{H}_{\lambda}]_{(\mathcal{K})}]_{|\Psi\rangle}\neq 0$ in the initial state, the QFI scales as $F_{\overline{\lambda}}\simeq N^{2(1+\mathcal{K})}/(\mathcal{K}!)^{2}$, establishing a direct and quantitative connection between the nilpotency index $\mathcal{K}$ and the degree of metrological enhancement. Through the QCRB, this scaling yields an RMSE $\Delta\overline{\lambda}\sim N^{-(1+\mathcal{K})}\mathcal{K}!$. This enhancement arises solely from the intrinsic non-commutativity between the auxiliary operator $\hat{H}_{g}$ and the parameter-encoding operator $\hat{H}_{\lambda}$, and requires no exotic quantum resources such as superpositions of causal orders that are often invoked in ICO-based schemes. Notably, the standard Heisenberg limit $\Delta\overline{\lambda}\sim N^{-1}$ emerges as a special case corresponding to $\mathcal{K}=0$, where no non-commutative structure is present.
Figure 2: (a) Logarithmic scaling of the coefficient of the leading-order QFI term ${N^{2(1+\mathcal{K})}}/{(\mathcal{K}!)^{2}}$, as a function of the number of operations NN, for fixed nilpotency indices $\mathcal{K}=1$, 4, and 6. Each curve represents the leading-order contribution from the series expansion of the total QFI in its respective encoding scheme, and exhibits a scaling $F_{\overline{\lambda}}^{(\mathcal{K})}\sim N^{2(1+\mathcal{K})}$, consistent with the structure of nested commutators. (b) Logarithmic scaling of the leading-order QFI coefficient as a function of the nilpotency index $\mathcal{K}$, for fixed $N=6$, 10, 16, and 20. For each given NN, a peak value $\mathcal{K}_{\mathrm{peak}}$ emerges (see Supplemental Material for its detailed expression [44]) which maximizes the contribution to the total QFI.
@@
However, when the $\mathcal{K}$-th order nested commutator becomes constant, for instance, in the case $\hat{H}_{\lambda}=\hat{P}=i(\hat{a}^{\dagger}-\hat{a})/\sqrt{2}$ and $\hat{H}_{g}=\hat{X}=(\hat{a}^{\dagger}+\hat{a})/\sqrt{2}$, where $\mathcal{K}=1$, the nested commutator $\hat{H}_{g},\hat{H}_{\lambda}]=i$, and hence $\text{Var}[[\hat{H}_{g},\hat{H}_{\lambda}]_{(\mathcal{K}=1)}]_{|\Psi\rangle}=0$, the RMSE scaling reduces to $\Delta\overline{\lambda}\sim N^{-1}$, recovering the standard Heisenberg limit. In this scenario, ICO-based schemes can unlock the enhanced scaling $\Delta\overline{\lambda}\sim N^{-(1+\mathcal{K})}\mathcal{K}!$ by introducing coherent superpositions of causal orders, which act as additional quantum resources to restore the contribution of the $\mathcal{K}$-th order nested commutator. Importantly, even under ICO, the metrological advantage is bounded by the underlying non-commutative structure between $\hat{H}_{g}$ and $\hat{H}_{\lambda}$ [44]. While ICO protocols cannot surpass the $\mathcal{K}$-dependent scaling limit, they can activate metrological enhancement that is otherwise inaccessible under definite causal order by coherently exploiting quantum control over operator sequences.
More broadly, an ICO-based scheme can be used to estimate either $\overline{\lambda}$ or $\overline{g}$. However, when the associated nilpotency index differ, i.e., $\mathcal{K}_{[\hat{H}_{g},\hat{H}_{\lambda}]}\neq\mathcal{K}_{[\hat{H}_{\lambda},\hat{H}_{g}]}$, the achievable precision for each parameter becomes intrinsically different. In such cases, the RMSE for each estimate is dictated by the nilpotency index corresponding to the relevant generator, depending on the asymmetric non-commutative structure of the encoding process [44].
Infinite nilpotency index $\mathcal{K}\to\infty$.—In this regime, by constructing a class of non-commutative encoding operators, we demonstrate that the infinite nilpotency can induce an exponential precision scaling of RMSE.
@@
demonstrating $\Delta\overline{\lambda}\simeq N^{-1}e^{-N}$ scaling driven by the underlying non-commutative structure of the encoding operators.
@@
Importantly, this enhancement is generic and does not require specially prepared probe states: exponential scaling holds for a variety of $|\Psi\rangle$ as long as $\text{Var}[i\hat{C}]_{|\Psi\rangle}\neq 0$, $\text{Var}[\hat{D}]_{|\Psi\rangle}\neq 0$, or $\text{Cov}[i\hat{C},\hat{D}]_{|\Psi\rangle}\neq 0$. This framework highlights the role of infinite nilpotency index $\mathcal{K}\to\infty$ as a necessary condition for exponential enhancement. Furthermore, it reveals the potential for systematically optimizing metrological performance by engineering the auxiliary operator $\hat{H}_{g}$ to tailor the non-commutative structure, for both finite and infinite $\mathcal{K}$.
@@
Examples.—We now propose several experimentally feasible protocols, implemented within continuous-variable (CV) systems, to realize the predicted precision enhancement. This focus stems from a fundamental limitation in discrete-variable (DV) systems: as rigorously proven in the Supplementary Material [44], the variance of the local generator is strictly bounded by the finite geometry of the generalized Bloch vector space for a $d$-dimensional system. This bound inherently precludes both the power-law ($N^{-(1+\mathcal{K})}$) and exponential ($N^{-1}e^{-N}$) scalings achievable in CV systems, limiting precision to at most oscillatory behavior with $N$. Consequently, ICO cannot be used to achieve super-Heisenberg scaling for DV quantum metrology, addressing an interesting open question in Ref. [32].
@@
(i) Finite nilpotency index $\mathcal{K}$.—Now we present a quantum metrological protocol that achieves enhanced scaling by exploiting non-commutative operations with a finite nilpotency index $\mathcal{K}=1$.
@@
The protocol involves a sequence of non-commutative operations generated by $\hat{H}_{g}=\hat{X}^{2}$ [45, 46], $\hat{H}_{\lambda}=\hat{P}$, implemented via the unitary evolution $\hat{U}_{\overline{s}}=\prod_{i=0}^{N}\hat{D}_{x_{i}}\prod_{j=0}^{N}\hat{S}_{s_{j}}$. Here $\hat{D}_{x_{i}}=e^{-ix_{i}\hat{P}}$ denote the momentum displacements and $\hat{S}_{s_{j}}=e^{-is_{j}\hat{X}^{2}}$ are single-mode shearing operations. The goal is to estimate the average displacement $\overline{x}=\sum_{i=0}^{N}x_{i}/N$, with the mean shearing strength $\overline{s}=\sum_{j=0}^{N}s_{j}/N$ treated as an auxiliary parameter.
@@
For this sequence, the nilpotency index between operators $\hat{X}^{2}$ and $\hat{P}$ is $\mathcal{K}=1$ while the relevant nested commutator is $\hat{H}_{g},\hat{H}_{\lambda}]_{(\mathcal{K}=1)}=2i\hat{X}$, yielding the local generator $\hat{h}_{\overline{x}}=N\hat{P}-2N^{2}\overline{s}\hat{X}$. The QFI takes the form: $F_{\overline{x}}=4N^{2}\text{Var}[\hat{P}]_{|\Psi\rangle}-4N^{3}\overline{s}\text{Cov}[\hat{X},\hat{P}]_{|\Psi\rangle}+4N^{4}\overline{s}^{2}\text{Var}[\hat{X}]_{|\Psi\rangle},$ where $\text{Cov}[\hat{X},\hat{P}]_{|\Psi\rangle}=\langle\hat{X}\hat{P}+\hat{P}\hat{X}\rangle-2\langle\hat{X}\rangle\langle\hat{P}\rangle$. For $\overline{s}\neq 0$, the $N^{4}$-scaling term dominates, leading via the QCRB to an RMSE scaling of $\Delta\overline{x}\sim N^{-2}$, surpassing the standard Heisenberg scaling. This enhancement originates purely from the non-commutativity between $\hat{X}^{2}$ and $\hat{P}$. However, when the relevant nested commutator is a constant (e.g., $\hat{H}_{g}=\hat{X}$, $\hat{H}_{\lambda}=\hat{P}$), ICO enables achieving the same super-Heisenberg scaling $N^{-2}$, a finding corroborated by prior works [32, 33] and consistent with our general result above.
@@
(ii) Infinite nilpotency index $\mathcal{K}\to\infty$.—We now turn to a quantum metrology protocol that achieves exponential enhancement in parameter estimation by a non-commutative process through sequential displacement $\hat{H}_{\lambda}=\hat{P}$ and squeezing operations $\hat{H}_{g}=(\hat{a}^{\dagger})^{2}+\hat{a}^{2}$.
@@
Figure 3: Logarithmic scaling of QFI $F_{\overline{x}}$ with respect to the number of operations NN for the non-commutative encoding protocol $\hat{U}_{x}=\prod_{j=1}^{N}\hat{D}_{x_{j}}\prod_{i=1}^{N}\hat{S}_{\xi_{i}}$. The probe state is a coherent state $|\Psi\rangle=\hat{D}(\alpha)|0\rangle$ with $\alpha=0.3$. Inset: The ratio between the classical Fisher information for a fixed quadrature measurement $Q_{\overline{x}}$ at $\theta=\pi/4$ and the QFI $F_{\overline{x}}$. The ratio consistently remains near unity (dashed line).

### v2

# Noncommutativity as a Universal Characterization for Enhanced Quantum Metrology

@@ Abstract
A central challenge in quantum metrology is to effectively harness quantum resources to surpass classical precision bounds. Although recent studies suggest that the indefinite causal order may enable sensitivities to attain the super-Heisenberg scaling, the physical origins of such enhancements remain elusive. Here, we introduce the nilpotency index $\mathcal{K}$, which quantifies the depth of noncommutativity between operators during the encoding process, can act as a fundamental parameter governing quantum-enhanced sensing. We show that a finite $\mathcal{K}$ yields an enhanced scaling of root-mean-square error as $N^{-(1+\mathcal{K})}$. Meanwhile, the requirement for indefinite causal order arises only when the nested commutators become constant. Remarkably, in the limit $\mathcal{K}\to\infty$, exponential precision scaling $N^{-1}e^{-N}$ is achievable. We propose experimentally feasible protocols implementing these mechanisms, providing a systematic pathway towards practical quantum-enhanced metrology.
@@
Introduction.—Quantum metrology represents one of the most promising near-term applications of quantum technologies [Maccone2004, Maccone2006, Romalis2007, PARIS2009, Cappellaro2017, Treutlein2018]. Its fundamental advantage stems from the ability of quantum resources—such as entanglement and coherence—to enhance measurement precision beyond the limits achievable with classical methods [Maccone2011, Zoller2024, Zelevinsky2024]. This is exemplified by the reduction in the root-mean-square error (RMSE) for parameter estimation from the standard quantum limit, scaling as $N^{-1/2}$, to the Heisenberg limit, scaling as $N^{-1}$, when utilizing $N$ resources [Geremia2007, Berry2020]. Such Heisenberg scaling has been demonstrated across diverse experimental platforms, achieved either by preparing $N$ entangled probes measured in parallel [Zeilinger2004, Silberberg2010, Pan2023] or by employing a single probe that evolves through and is measured after $N$ repetitions [Fig. 1(a)] [Pryde2007, Guo2018, Guo2021].
@@
To further push the sensitivity limit in measurement precision, recent processes involving indefinite causal order (ICO), where the sequence of fundamental encoding operations lacks definite temporal sequences, have been incorporated into quantum metrology [Walther2024, Brukner2014, Brukner2012, Rafa2023, Yang2023, Alastair2024]. A paradigmatic example is the quantum SWITCH [Fig. 1(b)], in which a control quantum system coherently dictates the order of two operations applied to a target system [Valiron2013, Walther2015, Walther2017, White2018, Giulio2024, Brukner2016, Brukner2016njp]. Significantly, using state-of-the-art photonic techniques, a precision scaling with super-Heisenberg limit $N^{-2}$ has been demonstrated [Liu2025, Guo2023].
@@
However, the fundamental mechanism by which ICO enhances measurement precision remains poorly understood, representing a key foundational question [Giulio2020, Guo2023, Zeng2024, Frey2019, Li2024, Ban2023, Delgado2023]. Understanding the essential resource underpinning this advantage holds significant promise for reducing experimental overhead while advancing the ultimate sensitivity limits. While it is often assumed that the coherence of causal order plays a central role, evidence suggests that simply increasing the number of orders does not necessarily lead to improved precision [footnote]. In existing protocols performing displacements generated by a given quadrature $\hat{X}$ and the conjugate quadrature $\hat{P}$ [Giulio2020, Guo2023, Liu2025] or two unitary gates $\hat{U}$ and $\hat{U}^{\dagger}$ for parameter encoding [Zeng2024], continuing to add similar paths brings little benefit and even makes the experiment harder to implement. Instead, the enhancement appears to stem from the noncommutativity between encoding operations in two paths, highlighting it as a more critical factor than only order multiplicity.
@@
Here, we establish a quantum metrology framework that exploits noncommutative sequences of encoding operations to enhance measurement precision. Central to our approach is the introduction of the noncommutativity depth $\mathcal{K}$, which is defined as the maximal integer yielding a nonzero nested commutator and serves as a fundamental quantity characterizing the noncommutative structure of the encoding operators. We rigorously demonstrate that $\mathcal{K}$ universally governs the scaling behavior of RMSE. For finite $\mathcal{K}$, the RMSE scales as $N^{-(1+\mathcal{K})}$ when the nested commutator remains nonconstant, enabling precision enhancement even without requiring the ICO scheme. However, when the $\mathcal{K}$-th order commutator becomes constant, ICO becomes necessary to unlock the optimal scaling. In the asymptotic limit $\mathcal{K}\to\infty$, we show that an exponential RMSE scaling $N^{-1}e^{-N}$ can be achieved, provided the nested commutators exhibit nonzero variance on the relevant initial state. Furthermore, we propose experimentally feasible protocols capable of achieving these enhanced scalings. Our findings offer a systematic approach to designing quantum control strategies for metrological optimization under realistic experimental conditions.
@@
General framework.—Generally, the precision of quantum metrology estimation governed by a unitary encoding $\hat{U}_{\lambda}=\text{exp}(-i\lambda\hat{H}_{\lambda})$ is fundamentally determined by the distinguishability between the neighboring output states $|\Psi_{\lambda}\rangle=\hat{U}_{\lambda}|\Psi\rangle$ and $|\Psi_{\lambda+\delta\lambda}\rangle=\hat{U}_{\lambda+\delta\lambda}|\Psi\rangle$. This distinguishability is quantified by the quantum Fisher information (QFI) $F_{\lambda}=4\text{Var}[\hat{h}_{\lambda}]_{|\Psi\rangle}$, where $\hat{h}_{\lambda}=i\hat{U}_{\lambda}^{\dagger}\partial_{\lambda}\hat{U}_{\lambda}$ denotes the local generator associated with $\lambda$, and the variance $\text{Var}[\cdots]_{|\Psi\rangle}$ is evaluated to the probe state $|\Psi\rangle$. The QFI sets a fundamental lower bound on the RMSE via the quantum Cramér-Rao bound (QCRB) $\Delta\lambda=1/\sqrt{\nu F_{\lambda}}$, where $\nu$ is the number of independent measurements. This bound is saturable in the asymptotic limit $\nu\to\infty$ through optimal choice of measurement and estimator [Cai2021, Caves1994, Augusto2009, Kok2010].
@@
Figure 1: The schematic illustrates representative quantum estimation schemes: (a) In the sequential protocol, a single probe state sequentially undergoes NN times of the parameter-encoding operator $\hat{H}_{\lambda}$. (b) In the ICO-based scheme, the probe evolves through a noncommutative encoding sequence whose order is coherently controlled by an auxiliary qubit. (c) In the general noncommutative encoding framework, the probe undergoes a sequence of $2N$ noncommutative encoding operations—NN times $\hat{H}_{g}$ and NN times $\hat{H}_{\lambda}$—executed in a definite order.
@@
We begin by introducing the general framework of the noncommutative metrology scheme [Fig. 1(c)]. The encoding unitary $\hat{U}_{\lambda}$ is constructed by sequentially applying two noncommutative operators, $\hat{H}_{\lambda}$ and $\hat{H}_{g}$ (i.e., $\hat{H}_{g},\hat{H}_{\lambda}]\neq 0$), and takes the form
@@
To quantify the depth of noncommutativity between $\hat{H}_{\lambda}$ and $\hat{H}_{g}$, we define the nilpotency index $\mathcal{K}\in\mathbb{N}$ via the adjoint action $\mathrm{ad}_{\hat{A}}(\hat{B})\equiv[\hat{A},\hat{B}$], such that
@@
We then define $\mathcal{K}$ as the maximal integer for which the nested commutator remains nonzero:
@@
That is, the action of $\mathrm{ad}_{\hat{H}_{g}}$ on $\hat{H}_{\lambda}$ terminates at order $\mathcal{K}+1$, $(\mathrm{ad}_{\hat{H}_{g}})^{\mathcal{K}+1}(\hat{H}_{\lambda})=0,$ which implies that $\mathrm{ad}_{\hat{H}_{g}}$ is nilpotent of index $\mathcal{K}+1$ on $\hat{H}_{\lambda}$. We note that the nilpotency index $\mathcal{K}$ is determined by the noncommutative algebraic structure and is independent of how many times the encoding operators are applied sequentially.
@@
Under the conditions of $\overline{g}\neq 0$ and a nonzero variance $\text{Var}[[\hat{H}_{g},\hat{H}_{\lambda}]_{(\mathcal{K})}]_{|\Psi\rangle}\neq 0$ in the initial state, the QFI scales as $F_{\overline{\lambda}}\simeq N^{2(1+\mathcal{K})}/(\mathcal{K}!)^{2}$, establishing a direct and quantitative connection between the nilpotency index $\mathcal{K}$ and the degree of metrological enhancement. Through the QCRB, this scaling yields an RMSE $\Delta\overline{\lambda}\sim N^{-(1+\mathcal{K})}\mathcal{K}!$. This enhancement arises solely from the intrinsic noncommutativity between the auxiliary operator $\hat{H}_{g}$ and the parameter-encoding operator $\hat{H}_{\lambda}$, and requires no exotic quantum resources such as superpositions of causal orders that are often invoked in ICO-based schemes. Notably, the standard Heisenberg limit $\Delta\overline{\lambda}\sim N^{-1}$ emerges as a special case corresponding to $\mathcal{K}=0$, where no noncommutative structure is present.
Figure 2: (a) Logarithmic scaling of the coefficient of the leading-order QFI term ${N^{2(1+\mathcal{K})}}/{(\mathcal{K}!)^{2}}$, as a function of the number of operations NN, for fixed nilpotency indices $\mathcal{K}=1$, 4, and 6. Each curve represents the leading-order contribution from the series expansion of the total QFI in its respective encoding scheme, and exhibits a scaling $F_{\overline{\lambda}}^{(\mathcal{K})}\sim N^{2(1+\mathcal{K})}$, consistent with the structure of nested commutators. (b) Logarithmic scaling of the leading-order QFI coefficient as a function of the nilpotency index $\mathcal{K}$, for fixed $N=6$, 10, 16, and 20. For each given NN, a peak value $\mathcal{K}_{\mathrm{peak}}$ emerges (see Supplemental Material for its detailed expression [sm]) which maximizes the contribution to the total QFI.
@@
However, when the $\mathcal{K}$-th order nested commutator becomes constant, for instance, in the case $\hat{H}_{\lambda}=\hat{P}=i(\hat{a}^{\dagger}-\hat{a})/\sqrt{2}$ and $\hat{H}_{g}=\hat{X}=(\hat{a}^{\dagger}+\hat{a})/\sqrt{2}$, where $\mathcal{K}=1$, the nested commutator $\hat{H}_{g},\hat{H}_{\lambda}]=i$, and hence $\text{Var}[[\hat{H}_{g},\hat{H}_{\lambda}]_{(\mathcal{K}=1)}]_{|\Psi\rangle}=0$, the RMSE scaling reduces to $\Delta\overline{\lambda}\sim N^{-1}$, recovering the standard Heisenberg limit. In this scenario, ICO-based schemes can unlock the enhanced scaling $\Delta\overline{\lambda}\sim N^{-(1+\mathcal{K})}\mathcal{K}!$ by introducing coherent superpositions of causal orders, which act as additional quantum resources to restore the contribution of the $\mathcal{K}$-th order nested commutator. Importantly, even under ICO, the metrological advantage is bounded by the underlying noncommutative structure between $\hat{H}_{g}$ and $\hat{H}_{\lambda}$ [sm]. While ICO protocols cannot surpass the $\mathcal{K}$-dependent scaling limit, they can activate metrological enhancement that is otherwise inaccessible under definite causal order by coherently exploiting quantum control over operator sequences.
More broadly, an ICO-based scheme can be used to estimate either $\overline{\lambda}$ or $\overline{g}$. However, when the associated nilpotency index differ, i.e., $\mathcal{K}_{[\hat{H}_{g},\hat{H}_{\lambda}]}\neq\mathcal{K}_{[\hat{H}_{\lambda},\hat{H}_{g}]}$, the achievable precision for each parameter becomes intrinsically different. In such cases, the RMSE for each estimate is dictated by the nilpotency index corresponding to the relevant generator, depending on the asymmetric noncommutative structure of the encoding process [sm].
Infinite nilpotency index $\mathcal{K}\to\infty$.—In this regime, by constructing a class of noncommutative encoding operators, we demonstrate that the infinite nilpotency can induce an exponential precision scaling of RMSE.
@@
demonstrating $\Delta\overline{\lambda}\simeq N^{-1}e^{-N}$ scaling driven by the underlying noncommutative structure of the encoding operators.
@@
Importantly, this enhancement is generic and does not require specially prepared probe states: exponential scaling holds for a variety of $|\Psi\rangle$ as long as $\text{Var}[i\hat{C}]_{|\Psi\rangle}\neq 0$, $\text{Var}[\hat{D}]_{|\Psi\rangle}\neq 0$, or $\text{Cov}[i\hat{C},\hat{D}]_{|\Psi\rangle}\neq 0$. This framework highlights the role of infinite nilpotency index $\mathcal{K}\to\infty$ as a necessary condition for exponential enhancement. Furthermore, it reveals the potential for systematically optimizing metrological performance by engineering the auxiliary operator $\hat{H}_{g}$ to tailor the noncommutative structure, for both finite and infinite $\mathcal{K}$.
@@
Examples.—We now propose several experimentally feasible protocols, implemented within continuous-variable (CV) systems, to realize the predicted precision enhancement. This focus stems from a fundamental limitation in discrete-variable (DV) systems: as rigorously proven in the Supplementary Material [sm], the variance of the local generator is strictly bounded by the finite geometry of the generalized Bloch vector space for a $d$-dimensional system. This bound inherently precludes both the power-law ($N^{-(1+\mathcal{K})}$) and exponential ($N^{-1}e^{-N}$) scalings achievable in CV systems, limiting precision to at most oscillatory behavior with $N$. Consequently, ICO cannot be used to achieve super-Heisenberg scaling for DV quantum metrology, addressing an interesting open question in Ref. [Guo2023].
@@
(i) Finite nilpotency index $\mathcal{K}$.—Now we present a quantum metrological protocol that achieves enhanced scaling by exploiting noncommutative operations with a finite nilpotency index $\mathcal{K}=1$.
@@
The protocol involves a sequence of noncommutative operations generated by $\hat{H}_{g}=\hat{X}^{2}$ [Filip2024, box2024], $\hat{H}_{\lambda}=\hat{P}$, implemented via the unitary evolution $\hat{U}_{\overline{s}}=\prod_{i=0}^{N}\hat{D}_{x_{i}}\prod_{j=0}^{N}\hat{S}_{s_{j}}$. Here $\hat{D}_{x_{i}}=e^{-ix_{i}\hat{P}}$ denote the momentum displacements and $\hat{S}_{s_{j}}=e^{-is_{j}\hat{X}^{2}}$ are single-mode shearing operations. The goal is to estimate the average displacement $\overline{x}=\sum_{i=0}^{N}x_{i}/N$, with the mean shearing strength $\overline{s}=\sum_{j=0}^{N}s_{j}/N$ treated as an auxiliary parameter.
@@
For this sequence, the nilpotency index between operators $\hat{X}^{2}$ and $\hat{P}$ is $\mathcal{K}=1$ while the relevant nested commutator is $\hat{H}_{g},\hat{H}_{\lambda}]_{(\mathcal{K}=1)}=2i\hat{X}$, yielding the local generator $\hat{h}_{\overline{x}}=N\hat{P}-2N^{2}\overline{s}\hat{X}$. The QFI takes the form: $F_{\overline{x}}=4N^{2}\text{Var}[\hat{P}]_{|\Psi\rangle}-4N^{3}\overline{s}\text{Cov}[\hat{X},\hat{P}]_{|\Psi\rangle}+4N^{4}\overline{s}^{2}\text{Var}[\hat{X}]_{|\Psi\rangle},$ where $\text{Cov}[\hat{X},\hat{P}]_{|\Psi\rangle}=\langle\hat{X}\hat{P}+\hat{P}\hat{X}\rangle-2\langle\hat{X}\rangle\langle\hat{P}\rangle$. For $\overline{s}\neq 0$, the $N^{4}$-scaling term dominates, leading via the QCRB to an RMSE scaling of $\Delta\overline{x}\sim N^{-2}$, surpassing the standard Heisenberg scaling. This enhancement originates purely from the noncommutativity between $\hat{X}^{2}$ and $\hat{P}$. However, when the relevant nested commutator is a constant (e.g., $\hat{H}_{g}=\hat{X}$, $\hat{H}_{\lambda}=\hat{P}$), ICO enables achieving the same super-Heisenberg scaling $N^{-2}$, a finding corroborated by prior works [Guo2023, Giulio2020] and consistent with our general result above.
@@
(ii) Infinite nilpotency index $\mathcal{K}\to\infty$.—We now turn to a quantum metrology protocol that achieves exponential enhancement in parameter estimation by a noncommutative process through sequential displacement $\hat{H}_{\lambda}=\hat{P}$ and squeezing operations $\hat{H}_{g}=(\hat{a}^{\dagger})^{2}+\hat{a}^{2}$.
@@
Figure 3: Logarithmic scaling of QFI $F_{\overline{x}}$ with respect to the number of operations NN for the noncommutative encoding protocol $\hat{U}_{x}=\prod_{j=1}^{N}\hat{D}_{x_{j}}\prod_{i=1}^{N}\hat{S}_{\xi_{i}}$. The probe state is a coherent state $|\Psi\rangle=\hat{D}(\alpha)|0\rangle$ with $\alpha=0.3$. Inset: The ratio between the classical Fisher information for a fixed quadrature measurement $Q_{\overline{x}}$ at $\theta=\pi/4$ and the QFI $F_{\overline{x}}$. The ratio consistently remains near unity (dashed line).